Show a complete solution to each problem
Adjusting antifreeze. Angela needs 20 quarts of 50% antifreeze solution in her radiator. She plans to obtain this by mixing some pure antifreeze with an appropiate amount of a 40% antifreeze solution. How many quarts of each should she use?

Solve each inequality. State the solution set in interval notation and sketch its graph.
1/6 - 1/3t > 0

Solve each inequality. State the solution set in interval notation and sketch its graph
-3/5 < 1/5 - 2/15w < -1/3

Solve each of the following problems by using an inequality.
Car Selling. Ronald wants to sell his car through a broker who charges a commission of 10% of the selling proce. Ronald still owes $11,025 on the car. Ronald must get enough to at least pay off the loan. What is the range of the selling price?

Answer by stanbon(75887)

Angela needs 20 quarts of 50% antifreeze solution in her radiator. She plans to obtain this by mixing some pure antifreeze with an appropiate amount of a 40% antifreeze solution. How many quarts of each should she use?
Let amount of 100% antifreeze be "x" quarts.
Then amount of 40% antifreeze is "20-x" quarts.
Antifreeze in the 100% part is 100%(x)=x quarts
Antifreeze in the 40% part is 40%(20-x)=(8-0.4x) quarts
EQUATION:
antifreeze + antifreeze = 50%(20) quarts
x+8-0.4x=10
0.6x=2
Multiply through by 10 to get:
6x=20
x=10/3 quarts (amounts of 100% antifreeze)
20-x=50/3 quarts (amount of 40% antifreeze)

Solve each inequality. State the solution set in interval notation and sketch its graph.
1/6 - 1/3t > 0
Multiply throu by 6 to get:
1-2t>0
Add 2t to both sides to ge:
2t<1
t<1/2
Interval solution: (-infinity,(1/2)
Solve each inequality. State the solution set in interval notation and sketch its graph
-3/5 < 1/5 - 2/15w < -1/3
Subtract (1/5) along the line to get:
-4/5 < (-2/15)w <-(1/3+1/5)
-4/5<-(2/15)w<-(8/15)
Multiply thru be 15 to get;
-12<-2w<-8
Divide thru by -2 to get:
4 Interval solution (4,6)

Solve each of the following problems by using an inequality.
Car Selling. Ronald wants to sell his car through a broker who charges a commission of 10% of the selling proce. Ronald still owes $11,025 on the car. Ronald must get enough to at least pay off the loan. What is the range of the selling price?
Let the selling price be "x"
INEQUALITY:
sell price - 10%(sell price)>=11,025
x-0.1x>=11,025
0.9x>=11,025
x>=$12,250
Cheers,
Stan H.